OpenStudy (smartballerina01):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 64t + 80 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

OpenStudy (.sammi.):

Well x-intercept means y=0, so \[f(t) = -16t^2 - 64t + 80 \\ \\ 0 = -16t^2 - 64t + 80 \] Can you factor this?

OpenStudy (smartballerina01):

yes I am factoring it right now

OpenStudy (smartballerina01):

Is it -16(t-1) (t+5) ?

OpenStudy (smartballerina01):

are the factors 1 and 5?

OpenStudy (smartballerina01):

I mean 1 and -5

OpenStudy (michele_laino):

correct!

OpenStudy (smartballerina01):

what does that mean about the x intercept of the funtion

OpenStudy (michele_laino):

the \(x-\)intercept, is the value of \(x\) such that \(f(x)=0\) in this case it is better to speak about \(t-\)intercept, such the function is \(f(t)\) and the \(t-\)intercepts are \(t=1,t=-5\)

OpenStudy (michele_laino):

it is easy, using the factorization above, to check, these conditions: \(f(1)=0,\;f(-5)=0\)

OpenStudy (smartballerina01):

f(1) = 0 and f(-5) = 0?

OpenStudy (michele_laino):

yes! For example if I substitute \(t=1\), I can write this: \[f\left( 1 \right) = - 16\left( {1 - 1} \right) \cdot \left( {1 + 5} \right) = - 16 \cdot 0 \cdot 6 = 0\] similarly, if I substitute \(t=-5\)

OpenStudy (phi):

also, the height f(t), in feet, of the sandbag if f(t) is zero , that means the height of the sandbag above the ground is 0 i.e. that is the bag hits the ground at that time.

OpenStudy (smartballerina01):

does it hit the ground at 5 or 1?

OpenStudy (michele_laino):

I think at \(t=1\), since, time has to be positive

OpenStudy (phi):

the f(-5) is 5 seconds before the sandbag was dropped. it is the "theoretical" time when you would have "launched" the bag up into the air to follow the parabola. Of course we don't do that... we start near the top by standing on a building

OpenStudy (michele_laino):

for part B hint: I add and subtract \(64\) so I get this: \[\begin{gathered} f\left( t \right) = - 16{t^2} - 64t - 64 + 64 + 80 = \hfill \\ \hfill \\ = \left( { - 16{t^2} - 64t - 64} \right) + 144 \hfill \\ \end{gathered} \] please continue

OpenStudy (smartballerina01):

are you completing the square?

OpenStudy (phi):

yes, that is showing how to complete the square

OpenStudy (smartballerina01):

I think i understand the question now. Thank you so much for your help